A room contains 54 chairs. The number of rows is 3 less that the number of chairs per row. Find the number of rows.

Answer by stanbon(75887)

A room contains 54 chairs. The number of rows is 3 less that the number of chairs per row. Find the number of rows.
----
# of rows = (# of chairs per row) - 3
r = c/r - 3
r = 54/r -3
r^2 + 3r -54 = 0
(r+9)(r-6) = 0
----
Positive solution:: # of rows = 6
-------
Cheers,
Stan H.
--------